# xTrimoPGLM: Unified 100B-Scale Pre-trained Transformer for Deciphering the Language of Protein

### v1

@@ Introduction
A significant limitation of xTrimoPGLM-100B is the high computational cost associated with the model, which presents a considerable barrier to their practical application. Fine-tuning without quantization requires at least four A100 80G GPUs, which may not be feasible for all users. To mitigate this, more advanced efficient compression technologies in terms of parameters or memory, such as quantization (?), kernel fusion (?) and other accelerate technologies (?, ?, ?, ?) could be applied. These methods might enable the training and deployment of larger models with reduced computational resources. In this area, there is a wealth of research exploring various methods. We have only validated a few of these methods so far. However, more research is needed to assess their effectiveness in real-world scenarios.
OOD Structure Prediction
@@
While xT-Fold has shown impressive both TM-score and inference speed in PLM-based models, a notable challenge persists in its OOD test performance (such as CASP proteins). This performance gap remains consistent, even as the pre-training model becomes more powerful. Though Figure1D illustrates that the model is gradually improving in compressing OOD sequences, it already diminishes the return. We speculate that protein sequences, unlike natural language, are less smooth in semantic space and more akin to a factual nature. Currently, it remains a formidable task for PLM-based methods to outperform those MSA-based (or retrieval-augmented), especially focusing on significant OOD test data. To bridge this gap, more abundant data resources, along with compute-optimal pre-training could be key. Another option is that integrating MSA modules with neural network retrievers (?, ?), might lead to better end-to-end optimization and faster inference speeds than traditional MSA methods. Such developments could enhance the model’s ability to generalize on OOD data while accelerating its processing capabilities.
Generative Protein Hallucinations
@@
In the preceding sections, we demonstrated that the xTrimoPGLM is capable of generating a diverse protein sequence. However, akin to most LLMs, our model also suffers from the issue of generating hallucinations. In the generation process, when the set sampling temperature is low, the model tends to generate fragments with a high repetition of amino acids. Although certain types of repetitive fragments (such as repeated Alanine) might be predicted with high-confidence structures, these fragments do not exist in nature. On the other hand, out of the 2,861 sequences we generated, approximately 21.7% of the sequences could not be confidently predicted by xT-Fold (pLDDT < 70), leaving us uncertain about whether these sequences can exist and fold stably. Even after SFT, only 17.8% to 88% of the generated sequences could find similar structures in the Protein Data Bank (PDB). To avoid or reduce the hallucination problem of LLMs, augmenting constraints (?, ?) during the training and sampling process can enhance the efficiency of the model’s generation.
Conclusion
@@
We explored unified understanding and generation pre-training with an extremely large-scale protein language model. Our experiments suggest that such scaling can extend to downstream tasks, including the key 3D structure prediction. We have further unlocked new possibilities in protein sequence design through supervised fine-tuning, paving the way for groundbreaking advancements in this field. Our work serves as a stepping stone for future research in the protein foundation model, and we hope it can facilitate further progress in a broader spectrum of protein-related applications.
Supplementary materials
1 Methods and Material
1.1 Backbone Framework: General Language Model (GLM)
@@
Current PLM frameworks are commonly categorized as either encoder-only, like ESM (?, ?), or decoder-only, such as ProGen (?, ?), exhibit limitations in addressing both task categories effectively due to their inherent inductive biases. Conceptually, the two types of tasks mirror the broader spectrum of protein sequence distributions (?, ?), suggesting the need for a unified model capable of encapsulating this diversity. To achieve this, encoder-decoder architectures like T5 (?) and non-causal decoder-only models like the General Language Model (GLM) (?, ?) are optimized through an auto-regressive generating and bidirectional input-processing objective, which emerge as promising candidates for this dual capability. However, the GLM, with its parameter efficiency, stands out as a more viable option compared to the T5, which requires significantly 2x larger parameters for similar efficacy. GLM is a transformer-based language model characterized by its unique training methodology. It employs autoregressive blank infilling while processing input text bi-directionally. This approach involves randomly blanking out continuous spans of tokens from the input and training the model to sequentially reconstruct these spans. This dual focus on autoencoding and autoregressive pre-training differentiates GLM from causal decoder-only language models, which rely solely on unidirectional attention.
1.2 Pre-Training Objectives
@@ 1.3 Pre-Training Strategy for xTrimoPGLM-100B
Masked Language Model Stage (400 Billion Pre-Trained Tokens). In this initial stage, the [MASK] token is employed for masking random tokens in the sequence, with these masked tokens comprising 15% of the total input. This stage, consuming about 400 billion tokens, is dedicated to advancing the model’s representation abilities.
Unified Training Stage (600 Billion Pre-Trained Tokens). Subsequently, the model undergoes training with a combined approach of MLM and GLM objectives, in a ratio of 20% MLM to 80% GLM. In this stage, the model processes an additional 600 billion tokens, aiming to further refine both its representational and generative capabilities.
@@
1.4 Empirical Analysis of Unified Training
This section presents an in-depth analysis of the feasibility of concurrently optimizing two distinct pre-training objectives in xTrimoPGLM. Unlike prior unified pre-training frameworks (?, ?, ?, ?, ?), which typically adopt similar formulations for diverse objectives, our approach extends these methodologies to a broader context. We critically examine whether models benefit from joint optimization of in-place token predictions (Masked Language Model) and next-token predictions (General Language Model). Central to our investigation are two pivotal questions: Objective Compatibility. Does the in-place token prediction objective be optimized with the next-token prediction approach simultaneously? This inquiry is essential to understand whether these objectives can be effectively integrated within a single training framework; Mutual Contribution. Can the in-place token prediction strategy enhance the performance of next-token prediction tasks, and does the reverse also hold true? This question addresses the potential synergistic effects of combining these two objectives in the training regime of xTrimoPGLM. Our exploration into these questions aims to shed light on the intricate dynamics of unified training models, particularly in the context of large-scale language models specialized for protein sequence analysis.
@@ 1.4 Empirical Analysis of Unified Training
We evaluate the compatibility of MLM (in-place token prediction) and GLM ([gMASK], next-token prediction) objectives. Each occupies 50% of the training batch time, alternating between them. Shifts in objectives occur at 100B and 200B token consumption milestones, facilitated by constant model parameters and architecture, requiring only adjustments in the attention mask. Validation losses (refer to Figure S8(a)(b)) indicate that despite their differing natures, both MLM and GLM objectives optimize simultaneously.
Furthermore, we investigate the impact of pre-training objectives on convergence speed. Models pre-trained on one objective adapt to another, training over an additional 50B tokens. Comparisons include: MLM-adapted GLM versus GLM trained from scratch and GLM-adapted MLM versus MLM trained from scratch. Our results (see Figure S8(c)(d)) show significantly faster convergence in adapted models compared to those trained from scratch. The MLM-adapted GLM matches the loss of the GLM from-scratch model with a 2.2$\times$ speedup (110B tokens). Similarly, the GLM-adapted MLM achieves a 2$\times$ speedup (100B tokens).
@@
1.5 The Training Stability of Unified Training
Training stability is a critical factor for the successful training of large language models (LLMs) at the 100B-scale (?, ?, ?, ?). Given a fixed computing budget, it is essential to balance efficiency and stability, particularly in relation to floating-point (FP) formats. Lower-precision FP formats, such as 16-bit precision (FP16), enhance computational efficiency but are vulnerable to overflow and underflow errors. These vulnerabilities can potentially lead to catastrophic collapses during training. xTrimoPGLM, drawing on the implementation strategies of GLM-130B (?), addresses many unstable training issues. Nonetheless, xTrimoPGLM-100B still experiences catastrophic training collapses during the transition from the first to the second stage, a challenge not present in smaller-scale models (10B-scale). Incorporating a fixed ratio of GLM loss into pre-training can trigger these collapses, even with a minimal 1% GLM loss ratio, as illustrated in Figure S9. To mitigate this issue, we propose the implementation of a smooth transition strategy.
@@ 1.5 The Training Stability of Unified Training
xTrimoPGLM-100B is trained on a cluster of 96 DGX-A100 GPU (8$\times$40G) servers in FP16 precision from January 18 to June 30, 2023. During this time, xTrimoPGLM-100B has consumed 1 trillion tokens from the dataset consisting of Uniref90 and ColabFoldDB. We adopt 3D parallel strategy with the 4-way tensor parallelism (?), 8-way pipeline parallelism (?), and 24-way data parallelism (?) based on DeepSpeed (?). The model owns 72 transformer layers, 80 attention heads, and 10,240 embedding dims with 31,744 feed-forward embedding dims using GeGLU (?). We adopt the Post-LN initialized with the DeepNorm (?). We follow the mixed-precision (?) strategy (Apex O2), i.e., FP16 for forwards and backward and FP32 for optimizer states and master weights, to reduce the GPU memory usage and improve training efficiency. We also adopt the Embedding Layer Gradient Shrink (EGS) strategy (?, ?) with $\alpha=0.1$ to stabilize the xTrimoPGLM-100B training. We warm up the batch size from 240 to 4224 over the first 2.5% samples. We use AdamW (?) as our optimizer with $\beta_{1}$ and $\beta_{2}$ set to 0.9 and 0.95, and a weight decay value of 0.1. We warm up the learning rate from $10^{-7}$ to $4\times 10^{-5}$ over the first 3.0% samples, then decay it by a 10 $\times$ cosine schedule to the minimum learning $4\times 10^{-6}$. We use a dropout rate of 0.1 and clip gradients using a clipping value of 1.0. Each sample contains a fixed sequence length of 2,048 (We concatenate all protein sequences with a separator into a single document, and sample protein sequences from this document in such a way that there is virtually no padding during pre-training.). To adapt the model to the different lengths of proteins in the downstream tasks, we adopt the mix-length pre-training strategy with four different context windows of 256, 512, 1,024, and 2,048. Taking, 512, for example, we concatenate four samples together to cater for the 2,048-sequence-length. The ratio of different context lengths is [#$[\#256:\#512:\#1,024:\#2,048=0.1:0.4:0.4:0.1]$. We implement the two-dimensional RoPE from its author blog https://kexue.fm/archives/8397 as our position embedding. For the tokenization of the protein data, we use the residue-level tokenizer which is adopted in several PLMs (?, ?, ?, ?). Except for the basic amino acid types, we add special tokens [MASK], [sMASK], and [gMASK] for model prediction. We also add special tokens <sop>, <eop>, <eos> for sequence separation (Cf. Table S8 for the full configurations).
1.7 xTrimoPGLM-100B Pre-Training Datasets
@@
The pre-training dataset of xTrimoPGLM-100B is curated from two extensive data repositories: Uniref90 (https://www.uniprot.org/help/downloads), the Uniref90 version preceding December 2022 is downloaded and ColabFoldDB (?)( https://colabfold.mmseqs.com). The initial contributions from Uniref90 and ColabFoldDB encompasses approximately 153M and 950M (210M representatives plus 740M members) entries, respectively.
Uniref, a cluster from UniProt, is broadly acknowledged as a high-quality protein dataset often utilized in pre-training PLMs such as ESM (?) and ProtTrans (?). UniRef90 clusters are generated from the UniRef100 seed sequences with a 90% sequence identity threshold using the MMseqs2 https://github.com/soedinglab/MMseqs2 algorithm, as depicted in Figure S11. The left panel illustrates the composition of the dataset used for pre-training the model. The right panel depicts the distribution of taxonomic categories of Uniref90, visualized as concentric circles representing the levels of superkingdom, kingdom, and phylum from innermost to outermost. The innermost circle represents four superkingdoms: Bacteria (67%), Archaea (3%), Eukarya (27%), and Viruses (1%), with 2% of sequences labeled as unclassified. The middle circle encompasses 17 classified kingdoms, including an unclassified bacteria category, denoted as “bacteria*”. The outermost circle denotes the phylum level, marking only those labels with counts over 200,000. In total, Uniref90 includes 273 known phyla. This comprehensive representation across multiple taxonomic levels demonstrates the rich biodiversity encapsulated within the Uniref90 dataset and affirms its value for wide-ranging biological investigations. Protein sequences that are published prior to January 1, 2023, are incorporated into the training set. Given its robustness and reliability, our training process also substantially prioritizes this dataset.
@@
ColabFoldDB is established through an amalgamation of various metagenomic databases including BFD https://bfd.mmseqs.com, MGnify (?), SMAG (eukaryotes) (?), MetaEuk (eukaryotes) (?), TOPAZ (eukaryotes) (?), MGV (DNA viruses) (?), GPD (bacteriophages) (?), and an updated version of the MetaClust (?) dataset. Built upon the foundation of UniProtKB, ColabFoldDB is substantially augmented with a large corpus of metagenomic sequences derived from diverse environmental niches. Metagenomic data introduces a new level of diversity to the database, encompassing numerous environmental niches ranging from the human gut to marine ecosystems. This offers unparalleled opportunities for the discovery of novel proteins. To comprehensively map the entirety of protein sources in the biological world, the pre-training dataset has been expanded by incorporating protein sequences sourced from ColabFoldDB in addition to those from the Uniref90 dataset.
@@
Training Set. The complete dataset in ColabFoldDB initially contained approximately  950M sequences. After initial deduplication and short-length filtering, which removed about  150M duplicate sequences, and further refinement by cross-referencing and deduplicating with Uniref90, we narrowed down the dataset to  780M unique sequences, ensuring diversity and representativeness for effective training. We conduct a composition analysis of each remaining sequence, excluding any that exhibit an individual amino acid composition exceeding 80% as this may indicate an anomaly or bias in the data. These steps leave us a more representative subset of around  200M sequences. Then, we combine the two refined datasets (i.e., Uniref90 and ColabFoldDB), yielding a collection of  940M unique sequences, equivalent to roughly 200 billion tokens, as depicted in Figure S11 (left). During training, to capitalize on the high-quality data, we assign a greater weight to the Uniref90 data, resulting in a ColabFoldDB sampling ratio of approximately 60%. This approach triples or quadruples the contribution of Uniref90 data, boosting our model’s fine-tuning capability with high-quality data.
@@
Pre-training Data Distribution. As shown in Figure S12, the bar charts represent the distribution of sequence lengths within the Uniref90 and ColabFoldDB datasets. In both datasets, sequences in the ’100-400’ length category predominate, followed by the ’50-100’ category. The ’0-50’ and ’400+’ categories contain significantly fewer sequences. Note the comparison between the distribution of Uniref90 and ColabFoldDB, indicating the variety of sequence lengths used for model training.
1.8 Downstream Tasks Datasets & Evaluation Metrics

### v2

@@ Discussion
Although xTrimoPGLM demonstrated impressive performance, it has limitations. A significant limitation of xTrimoPGLM-100B is the high computational cost associated with the model, which presents a considerable barrier to their practical application. Fine-tuning without quantization requires at least four A100 80G GPUs, which may not be feasible for all users. To mitigate this, more advanced efficient compression technologies in terms of parameters or memory, such as quantization  (?), kernel fusion  (?) and other accelerate technologies  (?, ?, ?, ?) could be applied. These methods might enable the training and deployment of larger models with reduced computational resources. In this area, there is a wealth of research exploring various methods. We have only validated a few of these methods so far. However, more research is needed to assess their effectiveness in real-world scenarios.
@@
Moreover, we observed diminishing returns in performance with increasing model parameters. Specifically, while xT-Fold has shown impressive both TM-score and inference speed in PLM-based models, a notable challenge persists in its OOD test performance (such as CASP proteins). This performance gap remains consistent, even as the pre-training model becomes more powerful. Though the model is gradually improving in compressing OOD sequences, it already diminishes the return (Figure 1D). We speculate that protein sequences, unlike natural language, are less smooth in semantic space and more akin to a factual nature. Currently, PLM-based methods still struggle to outperform MSA-based or retrieval-augmented approaches, especially when dealing with significant out-of-distribution (OOD) test data. To bridge this gap, leveraging more abundant data sources—such as sequences, structures, and functional descriptions  (?)—and expanding data modalities to include proteins, DNA, and RNA  (?), along with implementing compute-optimal pre-training strategies  (?), could be crucial. Therefore, we advocate for proportional scaling of both data and models and emphasize the importance of exploring their efficient frontier further. Another option is that integrating MSA modules with neural network retrievers  (?, ?, ?), could lead to better end-to-end optimization and faster inference speeds than traditional MSA methods. Such developments could enhance the model’s ability to generalize on OOD data while accelerating its processing capabilities.
@@
Similar to most large language models, xTrimoPGLM also experiences the issue of generating hallucinations. During the generation process, when the sampling temperature is set low, the model tends to produce fragments with a high repetition of amino acids. Although certain types of repetitive fragments (such as repeated Alanine) might be predicted with high-confidence structures, these fragments do not exist in nature. On the other hand, out of the sequences we generated, approximately 20% of the sequences could not be confidently predicted by xT-Fold (pLDDT < 70), leaving us uncertain whether these sequences can exist and fold stably. Even after SFT, only 17.8% to 88% of the generated sequences could find similar structures in the Protein Data Bank (PDB). To avoid or reduce the hallucination problem of LLMs, augmenting constraints  (?) during the training and sampling process can improve the efficiency of the model generation.
@@
To summary, we explored unified understanding and generation pre-training with an extremely large-scale protein language model. Our experiments suggest that such scaling can extend to downstream tasks, including the key 3D structure prediction. We have further unlocked new possibilities in protein sequence design through supervised fine-tuning, paving the way for groundbreaking advancements in this field. Our work serves as a stepping stone for future research in the protein foundation model, and we hope it can facilitate further progress in a broader spectrum of protein-related applications.
Acknowledgments
We extend our gratitude to Zhipu.AI for providing the computing resources. This work has been supported by the NSFC for Distinguished Young Scholar 62425601 and 62276148, New Cornerstone Science Foundation through the XPLORER PRIZE and research found from Zhipu.
Author contributions
B.C conceived the method, implemented and trained the model, investigated the scaling law, and drafted the manuscript. X.C. led the project, managed workflows, prepared pre-trained datasets, trained scaling law and xT-Fold, and refined the manuscript. P.L and Z.B analyzed the generated protein sequences. Y.G, J.G, S.L, and B.W evaluated the model on the protein understanding benchmark. X.T and X.Z helped to develop the xT-Fold. A.Z and C.L helped implement the model training framework. L.S and J.T served as the corresponding authors and played a crucial role in integrating all resources, with L.S serving as an intermittent advisor throughout. All the authors read and approved the manuscript.
Competing interests
The authors declare no competing interests.
Methods
Backbone Framework: General Language Model (GLM)
@@ Methods
Current PLM frameworks are commonly categorized as either encoder-only, like ESM  (?, ?), or decoder-only, such as PROGEN  (?, ?), exhibit limitations in addressing both task categories effectively due to their inherent inductive biases. Conceptually, the two types of tasks mirror the broader spectrum of protein sequence distributions  (?), suggesting the need for a unified model capable of encapsulating this diversity. To achieve this, encoder-decoder architectures like T5  (?) and non-causal decoder-only models like the General Language Model (GLM)  (?, ?) are optimized through an auto-regressive generating and bidirectional input-processing objective, which emerge as promising candidates for this dual capability. However, the GLM, with its parameter efficiency, stands out as a more viable option compared to the T5, which requires significantly 2x larger parameters for similar efficacy. GLM is a transformer-based language model characterized by its unique training methodology. It employs autoregressive blank infilling while processing input text bi-directionally. This approach involves randomly blanking out continuous spans of tokens from the input and training the model to sequentially reconstruct these spans. This dual focus on autoencoding and autoregressive pre-training differentiates GLM from causal decoder-only language models, which rely solely on unidirectional attention.
Pre-Training Objectives
@@
Masked Language Model Stage (400 Billion Pre-Trained Tokens). In this initial stage, the [MASK] token is employed for masking random tokens in the sequence, with these masked tokens comprising 15% of the total input. This stage, consuming about 400 billion tokens, is dedicated to advancing the model’s representation abilities. At this stage, the $\alpha$ for scaling the GLM loss is set to 0, focusing solely on minimizing the MLM loss to enhance the model’s understanding capabilities.
Unified Training Stage (600 Billion Pre-Trained Tokens). Subsequently, the model undergoes training with a combined approach of MLM and GLM objectives, in a ratio of 20% MLM to 80% GLM. In this stage, the model processes an additional 600 billion tokens, aiming to further refine both its representational and generative capabilities. Here, the $\alpha$ is set to 1, allowing both objectives to equally contribute to the overall loss function on each training instance, with the GLM objective appearing four times more frequently than the MLM objective.
@@
Empirical Analysis of Unified Training
This section presents an in-depth analysis of the feasibility of concurrently optimizing two distinct pre-training objectives in xTrimoPGLM. Unlike prior unified pre-training frameworks  (?, ?), which typically adopt similar formulations for diverse objectives, our approach extends these methodologies to a broader context. We critically examine whether models benefit from joint optimization of in-place token predictions (Masked Language Model) and next-token predictions (General Language Model). Central to our investigation are two pivotal questions: Objective Compatibility. Does the in-place token prediction objective be optimized with the next-token prediction approach simultaneously? This inquiry is essential to understand whether these objectives can be effectively integrated within a single training framework; Mutual Contribution. Can the in-place token prediction strategy enhance the performance of next-token prediction tasks, and does the reverse also hold true? This question addresses the potential synergistic effects of combining these two objectives in the training regime of xTrimoPGLM. Our exploration into these questions aims to shed light on the intricate dynamics of unified training models, particularly in the context of large-scale language models specialized for protein sequence analysis.
@@
We evaluate the compatibility of MLM (in-place token prediction) and GLM ([gMASK], next-token prediction) objectives. Each occupies 50% of the training batch time, alternating between them. Shifts in objectives occur at 100B and 200B token consumption milestones, facilitated by constant model parameters and architecture, requiring only adjustments in the attention mask. Validation losses indicate that despite their differing natures, both MLM and GLM objectives optimize simultaneously (Supplementary Figure 13(a)(b)).
Furthermore, we investigate the impact of pre-training objectives on convergence speed. Models pre-trained on one objective adapt to another, training over an additional 50B tokens. Comparisons include: MLM-adapted GLM versus GLM trained from scratch and GLM-adapted MLM versus MLM trained from scratch. Our results show significantly faster convergence in adapted models compared to those trained from scratch (Supplementary Figure 13(c)(d)). The MLM-adapted GLM matches the loss of the GLM from-scratch model with a 2.2$\times$ speedup (110B tokens). Similarly, the GLM-adapted MLM achieves a 2$\times$ speedup (100B tokens).
@@
The Training Stability of Unified Training
Training stability is a critical factor for the successful training of large language models (LLMs) at the 100B-scale  (?, ?, ?). Given a fixed computing budget, it is essential to balance efficiency and stability, particularly in relation to floating-point (FP) formats. Lower-precision FP formats, such as 16-bit precision (FP16), enhance computational efficiency but are vulnerable to overflow and underflow errors. These vulnerabilities can potentially lead to catastrophic collapses during training. xTrimoPGLM, drawing on the implementation strategies of GLM-130B  (?), addresses many unstable training issues. Nonetheless, xTrimoPGLM-100B still experiences catastrophic training collapses during the transition from the first to the second stage, a challenge not present in smaller-scale models (10B-scale). Incorporating a fixed ratio of GLM loss into pre-training can trigger these collapses, even with a minimal 1% GLM loss ratio (Supplementary Figure 12). To mitigate this issue, we propose the implementation of a smooth transition strategy.
@@
xTrimoPGLM-100B is trained on a cluster of 96 DGX-A100 GPU (8$\times$40G) servers in FP16 precision from January 18 to June 30, 2023. During this time, xTrimoPGLM-100B has consumed 1 trillion tokens from the dataset consisting of Uniref90 and ColabFoldDB. We adopt 3D parallel strategy with the 4-way tensor parallelism  (?), 8-way pipeline parallelism  (?), and 24-way data parallelism based on DeepSpeed  (?). The model owns 72 transformer layers, 80 attention heads, and 10,240 embedding dims with 31,744 feed-forward embedding dims using GeGLU  (?). We adopt the Post-LN initialized with the DeepNorm  (?). We follow the mixed-precision strategy (Apex O2), i.e., FP16 for forwards and backward and FP32 for optimizer states and master weights, to reduce the GPU memory usage and improve training efficiency. We also adopt the Embedding Layer Gradient Shrink (EGS) strategy  (?) with $\alpha=0.1$ to stabilize the xTrimoPGLM-100B training. We warm up the batch size from 240 to 4224 over the first 2.5% samples. We use AdamW  (?) as our optimizer with $\beta_{1}$ and $\beta_{2}$ set to 0.9 and 0.95, and a weight decay value of 0.1. We warm up the learning rate from $10^{-7}$ to $4\times 10^{-5}$ over the first 3.0% samples, then decay it by a 10 $\times$ cosine schedule to the minimum learning $4\times 10^{-6}$. We use a dropout rate of 0.1 and clip gradients using a clipping value of 1.0. Each sample contains a fixed sequence length of 2,048 (We concatenate all protein sequences with a separator into a single document, and sample protein sequences from this document in such a way that there is virtually no padding during pre-training.). To adapt the model to the different lengths of proteins in the downstream tasks, we adopt the mix-length pre-training strategy with four different context windows of 256, 512, 1,024, and 2,048. Taking, 512, for example, we concatenate four samples together to cater for the 2,048-sequence-length. The ratio of different context lengths is [#$[\#256:\#512:\#1,024:\#2,048=0.1:0.4:0.4:0.1]$. We implement the two-dimensional RoPE from its author blog https://kexue.fm/archives/8397 as our position embedding. For the tokenization of the protein data, we use the residue-level tokenizer which is adopted in several PLMs  (?, ?). Except for the basic amino acid types, we add special tokens [MASK], [sMASK], and [gMASK] for model prediction. We also add special tokens <sop>, <eop>, <eos> for sequence separation. (Cf. Table S12 for the full configurations).
Pre-Training Datasets
@@
The pre-training dataset of xTrimoPGLM-100B is curated from two extensive data repositories: Uniref90 (https://www.uniprot.org/help/downloads), the Uniref90 version preceding December 2022 is downloaded and ColabFoldDB  (?)( https://colabfold.mmseqs.com). The initial contributions from Uniref90 and ColabFoldDB encompass approximately 153M and 950M (210M representatives plus 740M members) entries, respectively.
@@
Uniref, a cluster from UniProt, is broadly acknowledged as a high-quality protein dataset often utilized in pre-training PLMs  (?, ?). UniRef90 clusters are generated from the UniRef100 seed sequences with a 90% sequence identity threshold using the MMseqs2( https://github.com/soedinglab/MMseqs2 algorithm). ColabFoldDB is established through an amalgamation of various metagenomic databases including BFD (https://bfd.mmseqs.com), MGnify  (?), SMAG (eukaryotes)  (?), MetaEuk (eukaryotes)  (?), TOPAZ (eukaryotes)  (?), MGV (DNA viruses)  (?), GPD (bacteriophages)  (?), and an updated version of the MetaClust  (?) dataset. Built upon the foundation of UniProtKB, ColabFoldDB is substantially augmented with a large corpus of metagenomic sequences derived from diverse environmental niches. Metagenomic data introduces a new level of diversity to the database, encompassing numerous environmental niches ranging from the human gut to marine ecosystems. This offers unparalleled opportunities for the discovery of novel proteins. To comprehensively map the entirety of protein sources in the biological world, the pre-training dataset has been expanded by incorporating protein sequences sourced from ColabFoldDB in addition to those from the Uniref90 dataset.
The left panel (Supplementary Figure 14) illustrates the composition of the dataset used for pre-training the model. The right panel depicts the distribution of taxonomic categories of Uniref90, visualized as concentric circles representing the levels of superkingdom, kingdom, and phylum from innermost to outermost. The innermost circle represents four superkingdoms: Bacteria (67%), Archaea (3%), Eukarya (27%), and Viruses (1%), with 2% of sequences labeled as unclassified. The middle circle encompasses 17 classified kingdoms, including an unclassified bacteria category, denoted as “bacteria*”. The outermost circle denotes the phylum level, marking only those labels with counts over 200,000. In total, Uniref90 includes 273 known phyla. This comprehensive representation across multiple taxonomic levels demonstrates the rich biodiversity encapsulated within the Uniref90 dataset and affirms its value for wide-ranging biological investigations. Protein sequences that are published prior to January 1, 2023, are incorporated into the training set. Given its robustness and reliability, our training process also substantially prioritizes this dataset.
@@
Training Set. The complete dataset in ColabFoldDB initially contained approximately  950M sequences. After initial deduplication and short-length filtering, which removed about 150M duplicate sequences, and further refinement by cross-referencing and deduplicating with Uniref90, we narrowed down the dataset to 780M unique sequences, ensuring diversity and representativeness for effective training. We conduct a composition analysis of each remaining sequence, excluding any that exhibit an individual amino acid composition exceeding 80% as this may indicate an anomaly or bias in the data. These steps leave us a more representative subset of around  200M sequences. Finally, the pre-trained dataset comprises approximately 939M protein sequences with 200B tokens. Specifically, the UniRef90 dataset contains around 156M protein sequences with 53B residue tokens. The ColabFoldDB cluster includes about 208M protein sequences with 38 tokens, and the ColabFoldDB member contains 575M sequences with 103B tokens. During training, to capitalize on the high-quality data, we assign a greater weight to the Uniref90 data, resulting in a ColabFoldDB sampling ratio of approximately 60%. This approach triples or quadruples the contribution of Uniref90 data, boosting our model’s fine-tuning capability with high-quality data.
@@
Pre-training Data Distribution. The bar charts (Supplementary Figure 15) represent the distribution of sequence lengths within the Uniref90 and ColabFoldDB datasets. In both datasets, sequences in the ’100-400’ length category predominate, followed by the ’50-100’ category. The ’0-50’ and ’400+’ categories contain significantly fewer sequences. Note the comparison between the distribution of Uniref90 and ColabFoldDB, indicating the variety of sequence lengths used for model training.
Data availability
All data used in this study are publicly available and the usages are illustrated in our methods. The pre-training dataset of xTrimoPGLM-100B is curated from two extensive data repositories: Uniref90 (https://www.uniprot.org/help/downloads), the Uniref90 version preceding December 2022 is downloaded and ColabFoldDB (https://colabfold.mmseqs.com). 18 downstream task datasets are all available online (https://huggingface.co/proteinglm). All structure prediction datasets are from AlphaFold Database (https://alphafold.ebi.ac.uk/download PDB database (https://www.rcsb.org/downloads) that released date is less than May 2020).
Code availability
Trained weight for the xTrimoPGLM model, and downstream datasets are available at https://huggingface.co/proteinglm.
Model training used DeepSpeed v0.6.1 https://github.com/microsoft/DeepSpeed. Data analysis used Python v3.8 https://www.python.org/, NumPy v1.16.4 https://github.com/numpy/numpy, SciPy v1.2.1 https://www.scipy.org/, seaborn v0.11.1 https://github.com/mwaskom/seaborn, Matplotlib v3.3.4 https://github.com/matplotlib/matplotlib, pandas v1.1.5 https://github.com/pandas-dev/pandas, TM-align v20190822 https://zhanglab.dcmb.med.umich.edu/TM-align/ was used for computing TM-scores. Structure visualizations were created in Pymol v2.3.0 https://github.com/schrodinger/pymol-open-source. Protein 3D structures were predicted using AlphaFold2 with the official implementations https://github.com/deepmind/alphafold.
Supplementary materials
1 Protein Understanding Benchmarks and Evaluations